OpenStudy (anonymous):

Using the table below, what is the change in enthalpy for the following reaction? CH4 (g) + 2O2 (g) arrow CO2 (g) + 2H2O (l) Substance ΔHf (kJ/mol) CO2 (g) -393.5 CO (g) -110.5 O2 (g) 0.0 O3 (g) 142.0 H2O (g) -241.8 H2O (l) -285.8 CH4 (g) -74.86 −890.2 kJ −604.4 kJ 560.4 kJ −802.2 kJ

OpenStudy (anonymous):

I'm not totally sure, i m also learning this stuff... so i think you should use this formula: \[\sum H= \sum nH_{f\ products}-\sum nH_{f\ reactans}\]

OpenStudy (anonymous):